A grindstone has moment of inertia 50m.k.s units .a constant torque is applied and the grindstone is found to have a speed of 150 r.p.m 10 seconds after starting from rest find the torque
1
Expert's answer
2020-04-08T10:54:28-0400
"\tau=I\alpha=I\frac{\Delta \omega}{\Delta t}"

"\tau=50\frac{150(2\pi)}{10\cdot 60}=78.5\ Nm"